# Ultrafast Extreme Events: Empirical Analysis of Mechanisms and Recovery in a Historical Perspective

### v1

@@ Abstract
To understand the emergence of Ultrafast Extreme Events (UEEs), the influence of algorithmic trading or high–frequency traders is of major interest as they make it extremely difficult to intervene and to stabilize financial markets. In an empirical analysis, we compare various characteristics of UEEs over different years for the US stock market to assess the possible non–stationarity of the effects. We show that liquidity plays a dominant role in the emergence of UEEs and find a general pattern in their dynamics. We also empirically investigate the after–effects in view of the recovery rate. We find common patterns for different years. We explain changes in the recovery rate by varying market sentiments for the different years.
@@ 1 Introduction
We present new results, extending the study in Ref. [21] to the years 2014 and 2021, and we compare the findings over the course of time with the time period 2007/2008, i.e. we study and compare the non–stationarity in the dynamics and recovery of UEEs from a period of human and machine trading to almost pure machine trading [28]. We want to see if there is a general pattern in the emergence and dynamics of UEEs due to the liquidity as its relevant driving mechanism. In Sec. 2 we introduce our data set. We define UEEs in Sec. 3 and show our results in Sec. 4. We conclude in Sec. 5.
@@ 4.1 Statistical Aspects of Ultrafast Extreme Events for 2014 and 2021
The 2–second criterion yields few additional UEEs of the order less than 12 %, suggesting that most are completed within 1.5 seconds. We find that UEEs rarely occur individually but are often clustered, i.e. when several UEEs are counted in one day, the following days often show a similar picture. For 2021, we notice clustering of UEEs as well. In 2014, the maximum number of UEEs found in one day is 33. In contrast to 2014, there is a significant increase in UEEs in January, more precisely in the period 25 January to 28 January 2021. The largest number of UEEs is counted on 28 January with 492 events, 444 of these occurred in the stock of American Airlines. This increase in UEEs can be explained by a Short Squeeze Event in January 2021 [32].
@@ 4.2 Mechanism for Ultrafast Extreme Events
Figure 9: Relative spread 400 spread events before and after a UEE begins, averaged across all UEEs for flash crashes (left) and flash spikes (right) in 2014 (top) and 2021 (bottom). Spread event zero indicates the start of a UEE according to Nanex, see Sec. 2.
The relative spread increases before the start of a UEE. Additionally, we notice large deviations between different stocks. To make a trend more visible, we average over the different spreads for a fixed spread event in Fig. 9. Here, the increase in liquidity before the start of a UEE is more clearly visible. This is consistent with the aforementioned observation that the largest returns in the quotes are associated with relatively small accumulated trading volumes. We emphasize that UEEs can occur even in usually liquid stocks when the liquidity drops momentarily.
@@ 4.3 Recovery of Stocks after Ultrafast Extreme Events
For each UEE in 2007/2008, 2014 and 2021 we show the recovery ratio for the first 50 to 100 trades after the UEE in Fig. 10 as contour plots. Clearly, for most of the flash crashes and flash spikes the prices at the end of a UEE stay on the extremum reached during the UEE. Importantly, a few flash crashes and flash spikes recover immediately after one trade. Furthermore, there are only a few flash crashes and flash spikes that generate an “aftershock”, i. e. the price continues to rise for flash spikes and fall for flash crashes after the UEE leading to a negative value for the recovery ratios. Moreover, most of the recovery ratios take values in the interval $0\leq\eta_{n}\leq 1$.
Figure 10: Contour plots of the recovery ratio (referred to as recovery rate in Ref. [21]) dependent on the trade number for flash crashes (left) and flash spikes (right) in 2007/2008 (top), in 2014 (middle) and 2021 (bottom). The top figure is taken from Ref. [21]
For the sake of clarity, we work out relative frequencies of events for a fixed $n$ where $\eta_{n}$ is larger than (or equal to) $0.8$ and where $\eta_{n}$ is smaller than (or equal to) $0.2$. In Fig. 11, these probabilities are displayed for $1\leq n\leq 100$.
Figure 11: Probability to have a recovery ratio of $\eta_{n}\geq 0.8$ (left) and $\eta_{n}\leq 0.2$ (right) for 2007/2008 (top), 2014 (middle) and 2021 (bottom). In the top figure, flash crashes are colored black, flash spikes gray. The top figure is taken from Ref. [21].
For $P(\eta_{n}\geq 0.8)$, i. e. for the probability for almost recovery, flash crashes and flash spikes are similar in 2014. The probability for almost recovery for flash crashes and flash spikes is faster in the years 2007/2008 for the very first trades. For $P(\eta_{n}\leq 0.2)$, i.e. the probability for almost no recovery, both flash crashes and flash spikes are very likely not to recover at the beginning. The probability of flash crashes is significantly larger than for flash spikes in 2007/2008. In 2021, the flash spikes do not recover as well as the flash crashes. The probability for almost no recovery afterwards decreases and remains largely constant for $50\leq n\leq 100$ for all years. Within the first 50 to 100 trades after a UEE it is determined whether a stock will recover almost completely or only incompletely from the shock caused by a UEE.
@@ 5 Conclusions
We carried out a large–scale empirical analysis and found that UEEs have universal characteristics in their formation and after–effects. To uncover this universality, we analyzed various years independently. As in a previous analysis of our group for 2007/2008 [21], we found that one return in the quotes during a UEE often almost fulfills the UEE criterion. Our analysis extends the study of Ref. [21]. To make a comparison of the results for the years 2014 and 2021 with the empirical analysis of Ref. [21] for 2007 feasible, we employed as a criterion of 2–seconds for the duration of UEEs. In addition, because the resolution of the 2014 and 2021 data is better, we also applied a 1.5–second criterion. Importantly, the number of UEEs is almost the same within less than 12 %. The statistical characteristics do not change much either. This leads to the important conclusion that the statistical properties of UEEs are fairly robust and universal when varying the criterion for their occurrence.
Most of the UEEs occur right after the pre– and after–market trading. In these time periods, the liquidity builds up or decreases again [33]. Both of these time periods are dominated by cancelations of limit orders. As the occurence of UEEs is a typical pattern for the years 2014 and 2021, it stands to reason that a large number of UEEs have similar underlying mechanisms. The liquidity status of the order book due to cancelations of limit orders is probably a crucial factor here. Gaps in the order book, which can be caused by cancelations of limit orders, can lead to larger return jumps in the midpoints [34]. From this point of view, it does not matter whether HFTs, other algorithmic traders or human traders are responsible for UEEs, as all types of traders can cancel limit orders and trigger larger price jumps through market orders. Presumably, these types of traders after the pre– and after–market trading play a minor role, as cancelations within this time periods seem to be a generic trading pattern. It is reasonable to assume that, even during a trading day, the underlying mechanism for many UEEs may be due to brief liquidity shortages, initiated by cancelations of limit orders and triggered by market orders. We collected evidence for this mechanism, but we do not exclude other mechanisms for UEEs. Further studies with a deeper look into the dynamics of the order book are called for.
@@
The larger returns in the quotes appear to be not necessarily directly responsible for the formation of UEEs, as these only appear after the start of the UEE. Rather, they might be important in the recovery process of a UEE. For all four years, we observe stocks with an almost full recovery after a UEE. The almost full recovery will be due to the state of the order book with less fluctuating liquidity after a UEE hinting at a more stable state of the order book. Even though their are similarities in the shape of the recovery ratio, the market sentiment causes non–stationary variations over the years. For instance, one reason for the difference between flash crashes and flash spikes in 2021 could have been the general stock market situation in that year. In spite of the short squeeze in 2021, the stock market year 2021 recovered from the COVID–19 pandemic [40, 41]. Traders may have tended to continue trading at higher price levels because it was assumed that the prices of stocks would continue to rise.

### v2

@@ Abstract
To understand the emergence of Ultrafast Extreme Events (UEEs), the influence of algorithmic trading or high–frequency traders is of major interest as they make it extremely difficult to intervene and to stabilize financial markets. In an empirical analysis, we compare various characteristics of UEEs over different years for the US stock market to assess the possible non–stationarity of the effects. We show that liquidity plays a dominant role in the emergence of UEEs and find a general pattern in their dynamics. We also empirically investigate the after–effects in view of the recovery rate. We find common patterns for different years. We explain changes in the recovery rate by varying market sentiments for the different years. Overall, our results hint at a certain degree of universal behavior.
@@ 1 Introduction
We present new results, extending the study in Ref. [21] to the years 2014 and 2021, and we compare the findings over the course of time with the time period 2007/2008, i.e. we study and compare the non–stationarity in the dynamics and recovery of UEEs from a period of human and machine trading to almost pure machine trading [28]. Here, “non-stationarity” refers to structural changes in market conditions across years, not to formal time-series stationarity. We want to see if there is a general pattern in the emergence and dynamics of UEEs due to the liquidity as its relevant driving mechanism. The consistency of the observed patterns shows a clear tendency towards universal behavior. As indicated in the title, our approach is purely empirical; we explore statistical properties and do not rely on a model interpretation which would always involve assumptions beyond empirical validation.
In Sec. 2 we introduce our data set. We define UEEs in Sec. 3 and show our results in Sec. 4. We conclude in Sec. 5.
@@ 4.1 Statistical Aspects of Ultrafast Extreme Events for 2014 and 2021
The 2–second criterion yields few additional UEEs of the order less than 12 %, suggesting that most are completed within 1.5 seconds. We find that UEEs rarely occur individually but are often clustered, i.e. when several UEEs are counted in one day, the following days often show a similar picture. For 2021, we notice clustering of UEEs as well. The clustering of UEEs can be interpreted in terms of short-term order-book instability. In market microstructure it is well established that temporary reductions in order-book depth and liquidity resilience tend to persist over short intervals, thereby increasing the likelihood of additional abrupt price movements within the same period [32, 33, 34]. Such short-lived liquidity shortages naturally generate clusters of extreme events. This microstructural interpretation is also consistent with broader frameworks of volatility clustering, including those discussed in the Fractal Market Hypothesis [35].
In 2014, the maximum number of UEEs found in one day is 33. In contrast to 2014, there is a significant increase in UEEs in January, more precisely in the period 25 January to 28 January 2021. The largest number of UEEs is counted on 28 January with 492 events, 444 of these occurred in the stock of American Airlines. This increase in UEEs can be explained by a Short Squeeze Event in January 2021 [36].
@@ 4.2 Mechanism for Ultrafast Extreme Events
Figure 9: Relative spread, 400 spread events before and after a UEE begins, averaged across all UEEs for flash crashes (left) and flash spikes (right) in 2014 (top) and 2021 (bottom). Spread event zero indicates the start of a UEE according to Nanex, see Sec. 2.
Figure 10: Standard deviations of relative spread, 400 spread events before and after a UEE begins, computed across all UEEs for flash crashes (left) and flash spikes (right) in 2014 (top) and 2021 (bottom). Spread event zero indicates the start of a UEE according to Nanex, see Sec. 2.
Figure 11: Adjusted Amihud illiquidity (AAI) of 400 spread events before and after a UEE begins, averaged across all UEEs for flash crashes (left) and flash spikes (right) in 2014 (top) and 2021 (bottom). Spread event zero indicates the start of a UEE according to Nanex, see Sec. 2.
The relative spread increases before the start of a UEE. Additionally, we notice large deviations between different stocks. To make a trend more visible, we average over the different spreads for a fixed spread event in Fig. 9. The corresponding standard deviations of the relative spread, computed across all UEEs for flash crashes and flash spikes, are depicted in Fig. 10. The averaged relative spread reveals a clear systematic trend. Here, the increase in liquidity before the start of a UEE is more clearly visible. This is consistent with the aforementioned observation that the largest returns in the quotes are associated with relatively small accumulated trading volumes. We emphasize that UEEs can occur even in usually liquid stocks when the liquidity drops momentarily. Furthermore, we study Amihud’s illiquidity ratio as one of the most widely used liquidity measures, as it combines price impact and trading activity into a single daily indicator [37]. Here, we introduce an adjusted form of Amihud’s illiquidity (AAI)
$l_{AAI}(t)=\frac{|r(t)|}{v_{\mathrm{avg}}(t)}\,$, (3)
where $|r(t)|$ denotes the absolute value of consecutive trade-to-trade returns and $v_{\mathrm{avg}}(t)$ is the average traded volume computed over the corresponding consecutive trades. However, applying the classical Amihud logic at millisecond resolution is not meaningful: at such ultrafast timescales, trade-to-trade returns are dominated by discreteness, bid–ask bounce, and mechanical order-book effects. Figure 11 clearly illustrates that the adjusted Amihud illiquidity measure, averaged across all UEEs, exhibits a fundamentally different time-scale behavior than the averaged relative spread in Fig. 9, further confirming that it cannot serve as a microstructural liquidity proxy at UEE timescales.
@@ 4.3 Recovery of Stocks after Ultrafast Extreme Events
For each UEE in 2007/2008, 2014 and 2021 we show the recovery ratio for the first 50 to 100 trades after the UEE in Figs. 12 and 13 as contour plots on a linear and logarithmic scale, respectively. The corresponding standard deviations of the recovery ratios, computed across all stocks for a fixed number of trades, are shown in Fig. 14. Clearly, for most of the flash crashes and flash spikes the prices at the end of a UEE stay on the extremum reached during the UEE. Importantly, a few flash crashes and flash spikes recover immediately after one trade. Furthermore, there are only a few flash crashes and flash spikes that generate an “aftershock”, i. e. the price continues to rise for flash spikes and fall for flash crashes after the UEE leading to a negative value for the recovery ratios. Moreover, most of the recovery ratios take values in the interval $0\leq\eta_{n}\leq 1$.
Figure 12: Contour plots of the recovery ratio (referred to as recovery rate in Ref. [21]) dependent on the trade number for flash crashes (left) and flash spikes (right) in 2007/2008 (top), in 2014 (middle) and 2021 (bottom) on a linear scale. The top figure is taken from Ref. [21]
Figure 13: Contour plots of the recovery ratio (referred to as recovery rate in Ref. [21]) dependent on the trade number for flash crashes (left) and flash spikes (right), in 2014 (top) and 2021 (bottom) on a logarithmic scale.
Figure 14: Standard deviations of the recovery ratio $\eta_{n}$ as a function of the number of trades, for flash crashes (left) and flash spikes (right) in 2014 (top) and 2021 (bottom).
For the sake of clarity, we work out relative frequencies of events for a fixed $n$ where $\eta_{n}$ is larger than (or equal to) $0.8$ and where $\eta_{n}$ is smaller than (or equal to) $0.2$. In Figs. 15 and 16, these relative frequencies are displayed for $1\leq n\leq 100$ on a linear and logarithmic scale, respectively.
Figure 15: Relative frequencies to have a recovery ratio of $\eta_{n}\geq 0.8$ (left) and $\eta_{n}\leq 0.2$ (right) for 2007/2008 (top), 2014 (middle) and 2021 (bottom) on a linear scale. In the top figure, flash crashes are colored black, flash spikes gray. The top figure is taken from Ref. [21].
Figure 16: Relative frequencies to have a recovery ratio of $\eta_{n}\geq 0.8$ (left) and $\eta_{n}\leq 0.2$ (right) for 2014 (top) and 2021 (bottom) on a logarithmic scale.
For $P(\eta_{n}\geq 0.8)$, i. e. for the relative frequencies for almost recovery, flash crashes and flash spikes are similar in 2014. The relative frequencies for almost recovery for flash crashes and flash spikes is faster in the years 2007/2008 for the very first trades. For $P(\eta_{n}\leq 0.2)$, i.e. the relative frequencies for almost no recovery, both flash crashes and flash spikes are very likely not to recover at the beginning. The relative frequencies of flash crashes is significantly larger than for flash spikes in 2007/2008. In 2021, the flash spikes do not recover as well as the flash crashes. The relative frequencies for almost no recovery afterwards decreases and remains largely constant for $50\leq n\leq 100$ for all years. Within the first 50 to 100 trades after a UEE it is determined whether a stock will recover almost completely or only incompletely from the shock caused by a UEE.
@@ 5 Conclusions
We carried out a large-scale empirical analysis and observed stable empirical patterns in the formation and after-effects of UEEs, which hint at a certatin degree of universal behavior. To uncover it, we analyzed various years independently. As in a previous analysis of our group for 2007/2008 [21], we found that one return in the quotes during a UEE often almost fulfills the UEE criterion. Our analysis extends the study of Ref. [21]. To make a comparison of the results for the years 2014 and 2021 with the empirical analysis of Ref. [21] for 2007 feasible, we employed as a criterion of 2–seconds for the duration of UEEs. In addition, because the resolution of the 2014 and 2021 data is better, we also applied a 1.5–second criterion. Importantly, the number of UEEs is almost the same within less than 12 %. The statistical characteristics do not change much either. This indicates that the statistical properties of UEEs are fairly stable with respect to variations in the event duration criterion.
Most of the UEEs occur right after the pre– and after–market trading. In these time periods, the liquidity builds up or decreases again [38]. Both of these time periods are dominated by cancelations of limit orders. As the occurence of UEEs is a typical pattern for the years 2014 and 2021, it stands to reason that a large number of UEEs have similar underlying mechanisms. The liquidity status of the order book due to cancelations of limit orders is probably a crucial factor here. Gaps in the order book, which can be caused by cancelations of limit orders, can lead to larger return jumps in the midpoints [39]. From this point of view, it does not matter whether HFTs, other algorithmic traders or human traders are responsible for UEEs, as all types of traders can cancel limit orders and trigger larger price jumps through market orders. Presumably, these types of traders after the pre– and after–market trading play a minor role, as cancelations within this time periods seem to be a generic trading pattern. It is reasonable to assume that, even during a trading day, the underlying mechanism for many UEEs may be due to brief liquidity shortages, initiated by cancelations of limit orders and triggered by market orders. We collected evidence for this mechanism, but we do not exclude other mechanisms for UEEs.
In line with this interpretation, UEEs occur predominantly in periods of low trading activity and sparse liquidity, where the reduced order flow limits the extent to which rapid reactive trading could reinforce the initial movement. The short and isolated structure of the events is therefore consistent with a liquidity-driven rather than a reaction-driven mechanism. Further studies with access to full order-book information would be valuable to explore this mechanism in more detail.
@@
The larger returns in the quotes appear to be not necessarily directly responsible for the formation of UEEs, as these only appear after the start of the UEE. Rather, they might be important for the recovery process of a UEE. For all four years, we observe stocks with an almost full recovery after a UEE. The almost full recovery will be due to the state of the order book with less fluctuating liquidity after a UEE hinting at a more stable state of the order book. Moreover, we find that the qualitative shape of the recovery patterns is remarkably robust across years with very different levels of market-wide volatility (2007/2008, 2014, 2021), indicating that the dimensionless recovery ratio is largely insensitive to broad volatility regimes. Even though there are similarities in the shape of the recovery ratio, the market sentiment causes non–stationary variations over the years. For instance, one reason for the difference between flash crashes and flash spikes in 2021 could have been the general stock market situation in that year. In spite of the short squeeze in 2021, the stock market year 2021 recovered from the COVID–19 pandemic [45, 46]. Traders may have tended to continue trading at higher price levels because it was assumed that the prices of stocks would continue to rise.